Ο ΤΙΤΛΟΣ ΜΟΥ

ΣΤΑΘΗΣ ΠΑΠΠΑΣ

March 26, 2025

1 ΕΝΟΤΗΤΑ

Η ΠΡΩΤΗ ΑΝΑΦΟΡΑ ΕΙΝΑΙ ΕΔΩ [?]

References

[1] David Ross-Pinnock and Paul G Maropoulos. Review of industrial temperature measurement technologies and research priorities for the thermal characterisation of the factories of the future. Proceedings of the Institution of Mechanical Engineers, Part B: Journal of Engineering Manufacture, 230(5):793–806, 2016. Publisher: IMECHE.